# Hard View Selection for Self-Supervised Learning

### v4

@@ 4 Main Results
Here, we discuss our main results on image classification, object detection, and segmentation tasks. All results are self-reproduced using the baseline code and available hyperparameters, and are averaged over 3 seeds (iBOT over 2 seeds). All hyperparameters are reported in Appendix G.
@@ 4.1 Evaluations on ImageNet
We report the top-1 validation accuracy on frozen features, as well as the k-NN classifier performance, in Table 1. For DINO, we additionally compare ResNet-50 (He et al., 2016) against the ViT-S/16 (Dosovitskiy et al., 2020) architecture. Our method compares favourably against all baselines with an increased performance of approximately $1\%$ on average, showing the benefit of sampling hard views. We highlight that our 300-epoch DINO+HVS model exhibits strong performance, closely rivaling the officially reported 800-epoch DINO performance of 77%, with a mere ${\sim}0.4\%$ performance gap. This achievement remains favorable, even when factoring in the 2$\times$ slowdown (DINO with 2 global, 8 local heads) induced by our method. Due to limited compute resources, we run the pretrainings for 100 and 300 epochs (200 epochs for SimCLR) and batch sizes of 512 (100 epoch) or 1024 (200 & 300 epoch trainings), respectively. This choice is in line with a strategy that favors the evaluation of a diverse and larger set of baselines over the evaluation of a less diverse and smaller set and underpins the broad applicability of HVS. We primarily ran our experiments with 8xNVIDIA GeForce RTX 2080 Ti nodes, with which the cheapest runs required ${\sim}3.5$ days for pretraining and linear evaluation and the most expensive runs required ${\sim}25$ days. Moreover, we emphasize that HVS is insensitive to the baseline hyperparameters and simply reusing these consistently resulted in improvements of the reported magnitudes across all experiments.
@@ 4.2 Transfer to Other Datasets and Tasks
We now demonstrate the transferability of features learned with HVS. For all transfer experiments, we use our 100-epoch ImageNet pretrained SimSiam ResNet-50 and DINO ViT-S/16 models.
@@
+ HVS ViT-S/16 95.13 98.65 81.27 88.23 92.07 93.60 49.03 54.16 74.13 87.91
Improvement +0.60 +0.12 +0.63 +0.33 +0.97 +0.40 +2.10 +0.19 +0.83 +0.41
Table 2: Models trained on ImageNet with HVS compare favourably against models trained without it when transferred to other datasets through finetuning (F.T.) or the linear evaluation protocol (Lin.).
Linear Evaluation and Finetuning
In Table 2, we apply both the linear evaluation (Lin.) and finetuning (F.T.) protocols to our models across a diverse set of datasets consisting of CIFAR10 Krizhevsky (2009), CIFAR100, Flowers102Nilsback & Zisserman (2008), Food101Bossard et al. (2014), and iNaturalist 2021 (iNaturalist 2021 competition dataset, ). Our results show that the improvements achieved by sampling hard views we observed so far also transfer to other datasets.
Method Arch. VOC Det. COCO Segm.
SimSiam RN50 78.46 45.16
+ HVS RN50 79.06 46.51
Impr. +0.60 +1.35
DINO RN50 80.86 50.35
+ HVS RN50 80.51 50.39
Impr. -0.36 +0.04
@@
Table 3: AP50 for object detection and instance segmentation.
Object Detection and Instance Segmentation
@@
For object detection, we use the VOC07+12 (Everingham et al., 2010) dataset using Faster R-CNN (Ren et al., 2015). We study instance segmentation on the COCO (Lin et al., 2014) dataset using Mask R-CNN (He et al., 2022). Table 3, where we report the AP50 performance, shows that the features learned with HVS also mostly transfer favourably to these tasks and outperform the SimSiam baseline with a 100-ep. pretraining. However, in case of DINO, we observe a slight performance gap on the detection task, which is the only instance in the entire paper where HVS did not outperform the baseline. This result warrants further investigation, as it was conducted with a single seed (due to limited compute resources). More details and performance results on this task are provided in Appendix F.1.
4.3 Time Complexity of HVS
Method Arch. Lin.
DINO ViT-S/16 74.06
+ HVS ViT-S/16 74.67
Impr. +0.61
SimSiam RN50 68.46
+ HVS RN50 68.98
Impr. +0.52
SimCLR RN50 64.31
+ HVS RN50 65.33
Impr. +1.02
Table 4: Accounting for computational overhead with 50% resolution (100 epochs).
The additional forward passes that HVS introduces for the selection phase increase the time complexity of the individual baseline methods. Several possible approaches exist to mitigate this overhead, one of which is to reduce the image resolution for the selection step (while still using the original resolution for learning). We measured the overhead factors under varying image resolutions for SimSiam and DINO and report them in Table 14 in the appendix. Given the 20% (SimSiam) to 30% (DINO) faster training times per batch when using a 50% image reduction, we evaluated this approach and observed that final performances are retained across all tested methods. Table 4 provides insights into the HVS’ performance when computational overhead is accounted for in the 50% reduction case. Overall, the results validate the effectiveness of HVS in reducing computational demands without compromising on performance. We refer to Section H in the appendix for a discussion on further optimizing HVS.
5 Empirical Analysis of Hard View Selection
@@
In this section, we discuss studies designed to shed light on the mechanisms underlying HVS. We address the following questions: 1) “Which can be observed that underlie the hard view selection?” 2) “Can a “manual” augmentation policy be inferred from these patterns?”, and 3) “What are the effects on empowering the adversary?”. For all experiments conducted here, we use our 100-epoch, ImageNet-pretrained SimSiam+HVS models with four sampled views.
5.1 Q1: Which Patterns Underlying the Hard View Selection Can be Observed?
@@
When visually studying examples and the views selected by HVS in Figures 2 and 5 (in the appendix), we notice that both geometric and appearance characteristics seem to be exploited, for instance, see the brightness difference between the views of the first two rows in Fig. 2. We also see a generally higher training loss (Fig. 6 in the appendix) indicative of an increased task difficulty.
Logging Augmentation Data
To assess these observations, we logged relevant hyperparameter data during SimSiam training with HVS. The logs include for each view the sampled geometric and appearance parameters from the data augmentation operations (such as the height/width of the crops or the brightness; see Section D in the appendix for more details), as well as the loss and whether the view was selected. As evaluated metrics we chose the Intersection over Union (IoU), Relative Distance (normalized distance of the center points views), color distortion distance (the Euclidean distance between all four color distortion parameters), and the individual color distortion parameters.
@@
Figure 3: Left: In over 40% of the cases, the adversarially selected view pair is also the one with the lowest IoU throughout SimSiam+HVS pretraining. We attribute the spike in the early phase to the random initialization of the embedding. Right: When using HVS (blue), a shift to smaller IoU values is visible when comparing to standard pretraining (orange). Both results are based on 3 seeds.
Importance of Augmentation Metrics
@@
Given 300k such samples, we then used fANOVA (Hutter et al., 2014) to determine how predictive these metrics are. This resulted in the metric with the highest predictive capacity on the loss to be the IoU, explaining 15% of the variance in performance, followed by brightness with 5% (for more details see Fig. 8 in the appendix). The importance of IoU in HVS is further underpinned by the following observation: the fraction of view pairs selected by HVS, which are also the pairs with the lowest IoU among all six pairs (N=4), is over 40% (random: ${\sim}16.7\%$) during training. Moreover, when using HVS, a shift to smaller IoU values can be observed when comparing against standard pretraining. We visualize both findings in Fig. 3.
Taking a Closer Look at the Intersection over Union
@@
We also examined the IoU value over the course of training in Fig. 4 (left). An observable pattern is that the IoU value with HVS (Fig. 4 (left) in blue) is smaller and varies more when compared against training without HVS (green). We believe this is due to the sample-wise and stateful nature of the adversarial selection as HVS chooses different IoU values between varying samples and model states. Lastly, we assessed the effect of the color augmentation on the pair selection. For this study, we sampled one set of color augmentations (as opposed to one for each view) per iteration and applied it to all views. Only after identifying the hardest pair, we apply sampled data augmentations to each view. As we show in Fig. 4 (right), the fraction of selected pairs that are also the hardest pairs slightly increases in this case. One possible explanation for this is that it reflects the non-negligible role of color variation between views (as shown previously with the importance analysis), where HVS is given less leverage to increase hardness through a static appearance and instead, depends more on leveraging the IoU. Another key observation is that HVS often chooses view pairs that incorporate zooming in and out or an increased distance between the views (see last row of Fig. 2). We believe the former relates to the analogy of local-to-global view correspondence described in SimCLR and the DINO papers.
5.2 Q2: Can a Manual Augmentation Policy be Inferred?
@@
Table 5: Top-1 lin. eval. accuracies for the manual IoU policy (averaged over 2 seeds).
@@
Since harder pretraining tasks seem beneficial according to observations made in Q1, a natural question arises: is it possible to mimic the adversarial selection with a manually scripted augmentation policy? Such a policy would replace HVS and lower the computational cost . Since the IoU plays an important role, below, we study several possible ways for constructing a simple manual augmentation policy based on IoU.
Deriving an Augmentation Policy
@@
We implemented the following rejection sampling algorithm in the augmentation pipeline: we linearly approximate the IoU values from Fig. 4 (left; in blue) with start and end values of 0.30 and 0.35 (ignoring the dip in the early phase). For each iteration, we then check if the pairs exceed the IoU value and if so, we reject the pair and re-sample a new pair. This ensures that only pairs are sampled that entail a minimum task difficulty (by means of a small enough IoU). We varied different hyperparameters, e.g., IoU start/end ranges, inverse schedules, and alternating between the IoU schedule and the standard augmentation . Training both SimSiam and DINO models for 100 epochs yielded performance decreases or insignificant improvements (see Table 5). These results indicate that developing a manual policy based on metrics in pixel space such as the IoU is non-trivial. Additionally, these results show that transferring such a policy from SimSiam to DINO does not work well, possibly due to additional variations in the augmentation pipeline such as multi-crop. In contrast, we believe that HVS is effective and transfers well since it 1) operates on a similarity level of latent embeddings that may be decorrelated from the pixel space and 2) has access to the current model state.
Figure 4: Left: The average IoU of view pairs selected by SimSiam+HVS (blue) compared against the default SimSiam training (green). Right: Using static color augmentation for all pairs before the selection increases the dependency on the IoU.
Assessing the Difficulty of Predicting the Pair Selected by HVS

### v6

@@ 5 Main Results
Here, we discuss our main results on image classification, object detection, and segmentation tasks. All results are self-reproduced using the original baseline code and hyperparameters (see Appendix Section A for details).
@@ 5.1 Evaluations on ImageNet
We report the top-1 validation accuracy on frozen features, as well as the k-NN classifier performance, in Table 1. For DINO, we additionally compare ResNet-50 (He et al., 2016) against the ViT-S/16 (Dosovitskiy et al., 2020) architecture. We point out that both methods, vanilla, and vanilla+HVP always receive the same number of data samples for training. Our method compares favorably against all baselines with an increased performance of approximately $1\%$ on average for 100 and 300 epoch pretraining, showing the benefit of sampling hard views.
Due to limited computing resources, we run the majority of pretrainings in this paper for 100 epochs and 300 epochs (200 epochs for SimCLR) and batch sizes of 512 (100 epoch) or 1024 (200 & 300 epoch trainings), respectively. This choice is in line with a strategy that favors the evaluation of a diverse and larger set of baselines over the evaluation of a less diverse and smaller set and underpins the broad applicability of HVP. We primarily ran our experiments with 8xNVIDIA GeForce RTX 2080 Ti nodes, with which the pretraining and linear evaluation duration ranged from ${\sim}3.5$ to ${\sim}25$ days. While HVP requires roughly 2x the training time of the DINO baseline (see Appendix J for further discussion on the time complexity of HVP), this investment translates directly into improved performance.
Rather than prioritizing efficiency, our focus was on achieving state-of-the-art results, highlighting HVP’s flexibility and robustness across training durations. To showcase this point, we explored the scalability of HVP with larger models and extended training schedules and achieved a new state-of-the-art result of 78.8% accuracy in linear evaluation, improving over the officially reported baseline of 78.2% on DINO ViT-B/16 (400 epochs). For k-NN classification, the same model similarly surpassed the DINO baseline by 0.85%, reaching 76.95% compared to 76.1%. These results demonstrate the scalability of our method to larger models and longer pretraining schedules. We emphasize that HVP is insensitive to the baseline hyperparameters and simply reusing the default ones consistently resulted in improvements in the reported magnitudes across all experiments.
@@ 5.2 Transfer to Other Datasets and Tasks
We now report the transferability of features learned with HVP. For all experiments here, we use our 100-epoch ImageNet pretrained iBOT and DINO ViT-S/16 models, respectively.
@@
+ HVP ViT-S/16 95.13 98.65 81.27 88.23 92.07 93.60 49.03 54.16 74.13 87.91
Impr. +0.60 +0.12 +0.63 +0.33 +0.97 +0.40 +2.10 +0.19 +0.83 +0.41
Table 2: HVP compares favorably against models trained without it when fine-tuned (F.T.) to or linearly evaluated (Lin.) on other datasets (averaged over 3 seeds; 100-ep. preraining).
5.2.1 Linear Evaluation and Finetuning
Method Arch. OD IS
100 300 100 300
iBOT ViT-S/16 66.13 66.80 63.10 63.63
+ HVP ViT-S/16 66.50 67.13 63.50 64.23
Impr. +0.37 +0.33 +0.40 +0.60
DINO ViT-S/16 65.90 66.60 62.83 63.63
+ HVP ViT-S/16 66.37 67.00 63.37 64.00
Impr. +0.47 +0.40 +0.53 +0.37
@@ 5.2.1 Linear Evaluation and Finetuning
Table 3: Object Detection (OD) and Instance Segmentation (IS) AP50 performance on COCO for 100/300 epoch pretraining.
In Table 2, we apply both the linear evaluation (Lin.) and finetuning (F.T.) protocols to our models across a diverse set of datasets consisting of CIFAR10 (Krizhevsky, 2009), CIFAR100, Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), and iNaturalist 2021 (iNaturalist 2021 competition dataset, ). Our results show that the improvements achieved by sampling hard views that we observed so far also transfer to other datasets.
5.2.2 Object Detection and Instance Segmentation
@@
For object detection and instance segmentation, we use the COCO (Lin et al., 2014) dataset with Cascade Mask R-CNN (Cai & Vasconcelos, 2019; He et al., 2017). Table 3, where we report the AP50 performance, shows that the features learned with HVP also transfer favorably to these tasks and outperform the iBOT and DINO baseline with a 100-ep. and 300-ep. pretraining. More details and performance results on this task are provided in Appendix H.1.
6 Empirical Analysis of HVP
@@
In this section, we discuss studies designed to shed light on the mechanisms underlying HVP. We address the following questions: 1) “Which pattern can be observed that underlies the hard view selection?” and 2) “What are the effects on empowering the adversary?”. For all experiments conducted here, we use our 100-epoch, ImageNet-pretrained SimSiam+HVP models with four sampled views. In Appendix F.1, we also analyze whether we can infer a “manual" augmentation policy from the following observed patterns.
6.1 Q1: Which Patterns Can be Observed with HVP?
@@
When visually studying examples and the views selected by HVP in Figures 1(b) and 5 (in the appendix), we notice that both geometric and appearance characteristics seem to be exploited, for instance, see the brightness difference between the views of the first two rows in Fig. 1(b). We also see a generally higher training loss (Fig. 6 in the appendix) indicative of an increased task difficulty.
@@
Figure 2: Left: In over 40% of the cases, the adversarially selected view pair has also the lowest Intersection over Union throughout SimSiam+HVP pretraining. We attribute the early spike to the random initialization of the embedding. Right: HVP (blue) shows a shift to smaller IoU values over standard pretraining (orange). Both results are based on 3 seeds.
6.1.1 Logging Augmentation Data
To assess these observations, we logged relevant hyperparameter data during SimSiam training with HVP. The logs include for each view the sampled geometric and appearance parameters from the data augmentation operations (such as the height/width of the crops or the brightness; see Section E in the appendix for more details), as well as the loss and whether the view was selected. As evaluated metrics, we chose the Intersection over Union (IoU), Relative Distance (normalized distance of the center points views), color distortion distance (the Euclidean distance between all four color distortion parameters), and the individual color distortion parameters.
6.1.2 Importance of Augmentation Metrics
@@
Given 300k such samples, we then used fANOVA (Hutter et al., 2014) to determine how predictive these metrics are. This resulted in the metric with the highest predictive capacity on the loss being the IoU, explaining 15% of the variance in performance, followed by brightness with 5% (for more details see Fig. 8 in the appendix). The importance of IoU in HVP is further underpinned by the following observation: the fraction of view pairs selected by HVP, which are also the pairs with the lowest IoU among all six pairs (N=4), is over 40% (random: ${\sim}16.7\%$) during training. Moreover, when using HVP, a shift to smaller IoU values can be observed when comparing against standard pretraining (see Fig. 2).
Figure 3: Left: The average IoU of view pairs selected by SimSiam+HVP (blue) compared against the default SimSiam training (green). Right: Using static color augmentation for all pairs before the selection increases the dependency on the IoU.
6.1.3 Taking a Closer Look at the Intersection over Union
We also examined the IoU value over the course of training in Fig. 3 (left). An observable pattern is that the IoU value with HVP (Fig. 3 (left) in green) is smaller and varies more when compared against training without HVP (blue). We believe this is due to the sample-wise and stateful nature of the adversarial selection as HVP chooses different IoU values between varying samples and model states.
@@ 6.1.3 Taking a Closer Look at the Intersection over Union
Lastly, we assessed the effect of the color augmentation on the pair selection. For this study, we sampled one set of color augmentations (as opposed to one for each view) per iteration and applied it to all views. We apply sampled data augmentations to each view only after identifying the hardest pair. As we show in Fig. 3 (right), the fraction of selected pairs that are also the hardest pairs slightly increases in this case. One possible explanation for this is that it reflects the non-negligible role of color variation between views (as shown previously with the importance analysis), where HVP is given less leverage to increase hardness through a static appearance and instead, depends more on leveraging the IoU. Another key observation is that HVP often chooses view pairs that incorporate zooming in and out or an increased distance between the views (see last row of Fig. 1(b)).
6.2 Q2: What are the Effects of Strong Adversaries?
It is well known that adversarial learning can suffer from algorithmic instability (Xing et al., 2021), e.g. by giving an adversary too much capacity. Here, we further explore the space of adversarial capacity for pretraining with hard views by adapting and varying HVP hyperparameters in order to gain a better understanding of their impact and robustness on discriminative learning. Additionally, we report further results on learning an adversary in Appendix G.1.
6.2.1 Robustness to Augmentation Hyperparameters
Figure 4: With HVP, SimSiam appears more robust to augmentation hyperparameter variation. We show this for RandomResizedCrop (left) and ColorJitter (right). For RRC, the values indicate the lower value of the sampling range and for CJ the intensity of the color cues. Results averaged over two seeds and SimSiam defaults are RRC=0.2 and CJ=0.5.
In assessing the robustness of HVP to augmentation hyperparameters, we conducted an analysis focusing on two primary augmentation operations: RandomResizedCrop (RRC) and ColorJitter (CJ). Our findings depicted in Fig. 4 suggest that HVP enhances the robustness of SSL methods, like SimSiam, against variations in these hyperparameters. Note, when varying one operation, either RRC or CJ, we maintained the other operation at its default configuration. In both settings, we observe less performance degradation with extreme augmentation values and overall smaller degradation rates for HVP. We believe that this robustness stems from hard view pretraining which inadvertently equips the model to handle stronger augmentations.
6.2.2 Increasing the Number of Views
In our initial experiments, we explored variations in the number of sampled views $N$ with SimSiam and HVP. As can be seen in Fig. 7 in the appendix, while $N=8$ views still outperform the baseline in terms of linear evaluation accuracy, it is slightly worse than using $N=4$ views (-0.05% for 100 epochs pretraining on linear evaluation). We interpret this result as the existence of a “sweet spot" in setting the number of views, where, in the limit, a higher number of views corresponds to approximating a powerful adversarial learner, capable of choosing very hard and unfavorable learning tasks that lead to model collapse and performance deterioration. We experimented with such an adversarial learner and report results in Appendix G.1.
7 Conclusion
We presented HVP, a new data augmentation and learning strategy for Self-Supervised Learning designed to challenge pretrained models with harder samples. This straightforward method allows pushing the effectiveness of the traditional random view generation in SSL. When combined with methods like DINO, SimSiam, iBOT, and SimCLR, HVP consistently showcased improvements of 1% on average in linear evaluation and a diverse set of transfer tasks. HVP achieved a new state-of-the-art result of 78.8% linear evaluation accuracy on DINO ViT-B/16, a 0.6% improvement over the previous baseline. This illustrates the scalability and effectiveness of our approach for larger pretraining settings. With growing models, there is an increasing demand for more data to effectively train them. Synthetic data generation offers a viable solution by enhancing the quantity and diversity of training data. Data augmentation techniques like HVP play a crucial role in this process by creating challenging views, which can serve as synthetic data. All in all, HVP holds promise for scenarios where one seeks to push absolute performance or explore making models less sensitive to hyperparameters, thereby strengthening them for various downstream applications.
8 Acknowledgements
Frank Hutter acknowledges the financial support of the Hector Foundation. We also acknowledge funding by the European Union (via ERC Consolidator Grant DeepLearning 2.0, grant no. 101045765). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Moreover, we acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828, by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG, as well as the Gauss Center for Supercomputing eV (www.gauss-centre.eu) for funding this project by providing computing time on the GCS supercomputer JUWELS at Jülich Supercomputing Center (JSC). We also acknowledge Robert Bosch GmbH for financial support and the computing time made available on the high-performance computer NHR@KIT Compute Cluster at the NHR Center NHR@KIT.
Appendix
Appendix A Reproducibility Statement
We provide complete code, environment installation instructions, hyperparameter settings and model checkpoints here: https://github.com/automl/hvp. For transparency, we outline all hyperparameters, data splits, and evaluation protocols in detail in Section I. Most experiments were run across multiple seeds, and we report average results to account for variability. Information regarding the required computational resources is discussed in Section J below.
Appendix B Examples Sampled by HVP
Figure 5: We depict row-wise ten example images from the ImageNet train set along with their sampled views with a finished, 100-epoch trained SimSiam ResNet50 model. Left: original image with the overlaid randomly sampled crops (colored dashed rectangles). Right: All views after applying resizing and appearance augmentations. The pair that is selected adversarially by HVP is highlighted in solid lines, eg. View 1 and View 4 in the first row.
Appendix C Training Loss
Figure 6: The training loss over 100 epochs. Comparing the DINO vanilla method with DINO + HVP. The spike and drop in the loss curve of DINO is caused by freezing the last layer in the first epoch which was proposed by the authors as a strategy to enhance downstream performance. For HVP we can only see a drop and no spike. We believe this is because HVP exposes the model to hard views from the beginning of training (i.e. the loss is immediately maximized).
Appendix D Effect of More Views on Linear Evaluation Performance
Appendix E Assessing the Importance of Metrics with fANOVA
To assess the importance of various metrics on the training loss, we apply fANOVA Hutter et al. (2014) on data that we logged during training with HVP. We used 300k samples that contain the following sampled parameters from the geometric and appearance data augmentation operations for each view: all random resized crop parameters (height and width of the original image, coordinates of crop corners and height and width of the crop), all Colorjitter (color distortion) strengths (brightness, contrast, saturation, hue), grayscale on/off, Gaussian blurring on/off, horizontal flip on/off, loss, and if the crop was selected or not. The metrics we chose are Intersection over Union (IoU), Relative Distance (sample-wise normalized distance of the center points of crop pairs), color distortion distance (the Euclidean distance between all four color distortion operation parameters, i.e. brightness, contrast, saturation, hue), and the individual color distortion parameters of the Colorjitter operation. As can be seen in Fig. 8, the metric with the highest predictive capacity on the loss is the IoU with an importance of 15.2% followed by brightness with 5.1%. The relative distance has an importance of 3.3%, the Colorjitter distance 2.3%, the contrast 1.6%, the saturation 1.4%, hue 0.6%, and all parameters jointly 1.7%.
Figure 8: Application of fANOVA Hutter et al. (2014) on logged training data to determine metrics with high predictive capacity on the training loss.
Appendix F Additional Empirical Analysis
F.1 Can a Manual Augm. Policy be Derived?
@@ Appendix F Additional Empirical Analysis
Table 4: Top-1 lin. eval. accuracies for the manual IoU policy.
@@
Since harder pretraining tasks seem beneficial according to observations made in Q1, a natural question arises: can we mimic the adversarial selection with a manually scripted augmentation policy? Such a policy would replace HVP and lower the computational cost. Since the IoU plays an important role, below, we study several ways to construct a simple manual augmentation policy based on IoU.
F.1.1 Deriving an Augmentation Policy
@@
We implemented the following rejection sampling algorithm in the augmentation pipeline: we linearly approximate the IoU values from Fig. 3 (left; in blue) with start and end values of 0.30 and 0.35 (ignoring the dip in the early phase). For each iteration, we then check if the pairs exceed the IoU value and if so, we reject the pair and re-sample a new pair. This ensures that only pairs are sampled that entail a minimum task difficulty (by means of a small enough IoU). We varied different hyperparameters, e.g., IoU start/end ranges, inverse schedules, and alternating between the IoU schedule and the standard augmentation . Training both SimSiam and DINO models for 100 epochs yielded performance drops or insignificant improvements (see Table 4). These results indicate that using a manual policy based on metrics in pixel space such as the IoU is non-trivial. Additionally, these results show that transferring such a policy from SimSiam to DINO does not work well, possibly due to additional variations in the augmentation pipeline such as multi-crop. In contrast, we believe that HVP is effective and transfers well since it 1) operates on a similarity level of latent embeddings that may be decorrelated from the pixel space and 2) has access to the current model state.
F.1.2 Assessing the Difficulty of Predicting the Pair Selected by HVP